### v2

# Robust Multimodal Learning via Cross-Modal Proxy Tokens

Abstract
Multimodal models often experience a significant performance drop when one or more modalities are missing during inference. To address this challenge, we propose a simple yet effective approach that enhances robustness to missing modalities while maintaining strong performance when all modalities are available. Our method introduces cross-modal proxy tokens (CMPTs), which approximate the class token of a missing modality by attending only to the tokens of the available modality. To efficiently learn the approximation for the missing modality via CMPTs with minimal computational overhead, we employ low-rank adapters in frozen unimodal encoders and jointly optimize an alignment loss with a task-specific loss. Extensive experiments on five multimodal datasets show that our method outperforms state-of-the-art baselines across various missing rates while achieving competitive results in complete-modality settings. Overall, our method offers a flexible and efficient solution for robust multimodal learning. The code and pretrained models will be released on GitHub.
1 Introduction
Figure 1: Our proposed approach, cross-modal proxy tokens (CMPTs), achieves state-of-the-art performance in both complete and missing modality scenarios. The radar plot shows F1-macro scores on the MM-IMDb dataset, with each axis representing the % availability of image and text modalities. We compare CMPTs against four recent methods [21, 29, 22, 20] using the same experimental setup as  [22, 20].
Figure 2: Illustration of our proposed method. We use pretrained unimodal encoders and adapt them to learn cross-modal proxy tokens (CMPTs) that can approximate missing modality class tokens. We learn the CMPTs and adaptation parameters by jointly optimizing alignment and task-specific objectives. Our approach can accommodate complete and missing modalities during both training and inference and enhances robustness across various missing modality scenarios.
Multimodal learning [35, 52] integrates information from multiple (diverse) input sources to improve the performance on downstream tasks. For example, examining a movie poster image along with a short text synopsis can provide a greater insight into the movie’s type and genre. In recent years, a number of methods have been proposed to effectively combine information from multiple modalities, including early-stage fusion [31], token-level fusion [49], channel exchange [47], and bottleneck mid-fusion [32]. Furthermore, specialized architectures have been proposed for specific multimodal applications, such as vision-language tasks [21, 23, 27], action recognition [51, 56], and segmentation [53, 40]. Most of these approaches assume that all modalities are available for every sample during both training and inference, yielding better performance than unimodal baselines. However, their performance drops sharply when one or more modalities get missing [39, 29].
In practical applications, input modalities can be missing during training and/or inference for various reasons; such as lack of synchronous data capture (limited paired data during training) and sensor failure or privacy concerns (modalities missing during inference). Ma et al. [29] observed that multimodal transformers such as ViLT [21] exhibit a drastic performance drop when one or more modalities are missing during inference as shown in Figure 1 (yellow curve). They developed an optimal fusion strategy to recover the performance loss of multimodal transformers in the missing modality scenarios (purple curve). Subsequently, Lee et al. [22] developed a prompt tuning technique (pink and blue curves), and Kim et al. [20] combined the prompt-based approach with Variance Covariance Invariance regularization (VICReg) (brown curve) to further improve the robustness of multimodal models to missing modalities. Achieving high performance when modalities are missing at varying rates remains a challenge, which we seek to address in this paper.
In this paper, we introduce a simple yet effective framework for multimodal learning that leverages pre-trained unimodal encoders to approximate missing modalities from the available ones. Specifically, we introduce cross-modal proxy tokens (CMPTs) that learn an approximation of the class tokens of the missing modality by attending on all the tokens of the available modality to handle different missing scenarios. As illustrated in Figure 2, our framework integrates pre-trained unimodal encoders into a multimodal learning setting while introducing CMPTs to bridge the performance gap in missing modality scenarios. To ensure efficient learning of CMPTs with minimal additional learnable parameters, we employ low-rank adapters within the frozen unimodal encoders and jointly optimize both an alignment loss and a task-specific loss. The radar plot in Figure 1 demonstrates the superior performance of our approach compared to five existing methods [21, 22, 20, 29] across six different complete and missing modality scenarios on the MM-IMDb dataset, where our CMPT-based model consistently outperforms all baselines. We conduct extensive experiments on five multimodal datasets, comparing our approach against existing state-of-the-art baselines. Results show that our framework achieves state-of-the-art performance in both missing modality (Section 4.3) and complete modality (Section 4.4) scenarios, while maintaining a minimal number of learnable parameters. Ablation studies further confirm that our method significantly improves performance even in severe missing-modality scenarios and provides performance enhancement across most of the classes, leading to an overall performance boost (Section 4.6).
Main contributions of our work are as follows.
  • We propose cross-modal proxy tokens (CMPTs) that learn an approximation of the missing modality from the available modality with minimal adaption of pre-train unimodal encoders via an alignment loss.
  • Our approach outperforms 12 recent baseline methods across four multimodal datasets in various missing-modality scenarios (Section 4.3).
  • Across five datasets, our method achieves on-par or superior performance compared to state-of-the-art techniques when all modalities are present during training and inference (Section 4.4).
  • We demonstrate that CMPTs generalize effectively to unseen missing rates (Section 4.5) and boost performance in most of the classes (Section 4.6).
2 Related Work
Recent works [29, 14, 39] show that performance can degrade significantly when modalities are missing, as multimodal models are not inherently robust to such scenarios. Several approaches have been proposed to address this issue. In this section, we primarily discuss works where the main focus is on making multimodal models robust to missing modalities. More broader related work on multimodal learning is discussed in Supplementary Section S1.
Robust model design is one of the approaches for handling missing modalities. Recently, Shin et al. [41] developed a robust framework utilizing modality masking and knowledge distillation. Wang et al. [44] proposed ShaSpec to impute missing features through modality-specific and modality-shared feature learning. Wang et al. [48] proposed TokenFusion, which dynamically replaces uninformative tokens to enhance robustness. Additionally, different fusion strategies [5, 10, 25] have been developed to improve robustness across various missing-modality scenarios. However, many of these approaches are task-specific, limiting their generalization to other multimodal tasks. In our work, we focus on reusing the pre-trained unimodal encoders, making it flexible for extension to different input modalities and multimodal tasks.
Robust training and model adaptation can enhance missing modality robustness. Recently Reza et al. [39] proposed a parameter-efficient adaptation framework for different multimodal tasks. Nezakati et al. [34] trained a single model to infer missing information by projecting available modalities. Ma et al. [28] introduced SMIL, a Bayesian meta-learning method for flexible handling of missing modalities during both training and testing. Ma et al. [29] proposed a multi-task optimization approach with a differentiable algorithm for optimal fusion strategy selection. Maheshwari et al. [30] developed a semi-supervised framework that leverages unlabeled data to improve robustness, while Wei et al. [50] used a teacher network to transfer multimodal information for improved missing modality robustness. These methods, however, either require extensive adaptation or specialized training procedures. In our approach, we employ lightweight adaptation of the unimodal encoders (using Rank-1 LoRA adapters) and a single extra loss term to explicitly learn cross-modal relationships. Hence, the overall training flow is mostly left unchanged.
Prompt tuning methods improve missing modality performance by incorporating learnable prompts across multiple layers of the model. Lee et al. [22] introduced missing-aware prompts that learns a separate set of prompts for each modality combination. Later, Jang et al. [16] simplified the approach, showing that a single set of prompts per modality can achieve similar performance. More recently, Dai et al. [6] proposed a multi-step adaptive prompt learning method through modality alignment, while Kim and Kim [20] introduced a read-only prompt tuning approach. Although effective, these methods require learning and storing a large number of prompts, typically 16 prompts [22, 16] per modality per layer, for each missing modality scenario. In contrast, our method uses a single cross-modal proxy token per modality to estimate the missing modality feature from the available modality, thereby reducing the complexity. We carry out extensive experimental comparisons with all the above prompt-based techniques to demonstrate superior performance across various missing modality scenarios. Furthermore, our approach is efficient, uses few learnable parameters, and is robust to missing modalities in both training and testing which we discuss in Section 4.
3 Proposed Approach
In this paper, we propose a new framework to train multimodal systems such that they retain their performance under different scenarios of missing modalities during inference. Furthermore, we seek to develop a simple and efficient framework that can be used for different multimodal tasks and datasets without any specialized (task-specific) changes. To achieve these objectives, we propose to leverage modality-specific (transformer-based) encoders that are pre-trained on large datasets and align the features across modalities using proxy tokens, as illustrated in Figure 2.
3.1 Multimodal embedding tokens
Let us consider a multimodal dataset containing two modalities =$\mathcal{M}=\{m_{1},m_{2}\}$. We denote the available dataset as $\mathcal{D}=\{\mathcal{D}_{c},\mathcal{D}_{m_{1}},\mathcal{D}_{m_{2}}\}$, which can be further divided into three subsets: $\mathcal{D}_{c}=\{x_{m_{1}},x_{m_{2}},y\}$ where both modalities are present with corresponding output labels $y$; $\mathcal{D}_{m_{1}}=\{x_{m_{1}},\varnothing,y\}$ where only modality $m_{1}$ $(x_{m_{1}})$ is available with corresponding labels $y$; $\mathcal{D}_{m_{2}}=\{\varnothing,x_{m_{2}},y\}$ where only modality $m_{2}$ $(x_{m_{2}})$ is available with corresponding labels $y$. Here, $\varnothing$ denotes a placeholder for the missing modality (e.g., empty string for text, zeros for image, audio, or video) following [22, 39].
For each input modality, we assume that a transformer-based encoder, pre-trained on some large dataset, is available. Specifically, each modality $m\in\mathcal{M}$ is processed through a pre-trained embedding layer, $\text{EmbeddingLayer}_{m}$, with parameters $\Theta_{e,m}$. The input $x_{m}$ for each modality is passed through its corresponding embedding layer to generate modality-specific tokens as
$\mathcal{T}_{m}=\text{EmbeddingLayer}_{m}(x_{m};\Theta_{e,m}),$ (1)
where $\mathcal{T}_{m}\in\mathbb{R}^{N\times d}$ denotes $N$ tokens, each of dimension $d$, for modality $m$. During training, we keep the pretrained parameters $\Theta_{e,m}$ frozen.
3.2 Cross-modal proxy tokens
To compensate for the missing modalities, we introduce cross-modal proxy tokens (CMPTs) that approximate the missing modality by attending only to the tokens of the available modality. For each modality $m\in\mathcal{M}$, we concatenate one CMPT ($\mathcal{T}_{\text{CMPT},m}\in\mathbb{R}^{1\times d}$) with the class token (CLS) ($\mathcal{T}_{\text{CLS},m}\in\mathbb{R}^{1\times d}$) and all the modality specific tokens $\mathcal{T}_{m}$ to generate the modified embedding tokens $\tilde{\mathcal{T}}_{m}\in\mathbb{R}^{(N+2)\times d}$ as
$\tilde{\mathcal{T}}_{m}=\text{Concat}(\{\mathcal{T}_{\text{CMPT},m},\mathcal{T}_{\text{CLS},m},\mathcal{T}_{m}\}).$ (2)
Note that the proposed CMPT serves as an extra token, similar to the class token [8, 7], but with a slightly orthogonal purpose. Class token attends to all the tokens of a given modality to aggregate information and learn a rich feature representation of that modality. In contrast, CMPT attends to all the tokens of a given modality to learn an approximation of the class token of the other modality (e.g., $\mathcal{T}_{\text{CMPT},m_{1}}$ approximates $\mathcal{T}_{\text{CLS},m_{2}}$ while attending to $\mathcal{T}_{m_{1}}$).
3.3 Learn CMPTs by alignment and adaptation
For each modality, we assume that a pretrained transformer encoder $f_{m}$ with parameters $\Theta_{f,m}$ is available. The encoder for each modality needs to be updated so that they can learn the corresponding CMPTs. To keep the number of trainable parameters minimal, we use low-rank adaptation (LoRA) [15] to learn a small number of parameters $\Delta_{f,m}$. The modified embedding tokens for modality $m$, $\tilde{\mathcal{T}}_{m}$, are then passed through the adapted encoder for modality $m$ to get the final output tokens as
$\hat{\mathcal{T}}_{m}=f_{m}(\tilde{\mathcal{T}}_{m};\Theta_{f,m},\Delta_{f,m}),$ (3)
where $\hat{\mathcal{T}}_{m}\in\mathbb{R}^{(N+2)\times d}$ represents the CMPT, CLS, and remaining tokens from the last encoder layer. Note that the pretrained encoder parameters $\Theta_{f,m}$ remain frozen while we learn $\Delta_{f,m}$.
We aim to train the CMPTs such that they act as a proxy of the class token of the missing modality at the inference time. In particular, we want the CMPT $\hat{\mathcal{T}}_{\text{CMPT},m_{1}}$ of modality $m_{1}$ to approximate the class token $\hat{\mathcal{T}}_{\text{CLS},m_{2}}$ of modality $m_{2}$, and vice versa. To achieve this, we use an alignment loss $\mathcal{L}_{\text{align}}$, defined as
$\mathcal{L}_{\text{align}}=\frac{1}{N_{\mathcal{D}_{c}}}\sum_{x\in\mathcal{D}_{c}}\begin{array}[]{l}\mathcal{L}_{\text{MSE}}(\hat{\mathcal{T}}_{\text{CMPT},{m_{1}}},\hat{\mathcal{T}}_{\text{CLS},{m_{2}}})\,+\\ \mathcal{L}_{\text{MSE}}(\hat{\mathcal{T}}_{\text{CMPT},{m_{2}}},\hat{\mathcal{T}}_{\text{CLS},{m_{1}}}),\end{array}$ (4)
where $\mathcal{L}_{\text{MSE}}$ denotes the mean squared error loss and ${N_{\mathcal{D}_{c}}}$ denotes the number of samples in the training data where both modalities are present.
3.4 Gating module and fusion
After computing the final output tokens $\hat{\mathcal{T}}_{m}$, we extract the class token ($\hat{\mathcal{T}}_{\text{CLS},m}$) and the CMPT ($\hat{\mathcal{T}}_{\text{CMPT},m}$) for both modalities and pass them through the gating module. If both modalities are available, the gating module returns the class tokens from each modality. If one modality is missing, the gating module substitutes the missing class token with the CMPT from the available modality. The returned tokens are then fused using additive fusion to produce the final fused feature token $\mathcal{T}\in\mathbb{R}^{1\times d}$ as
$\mathcal{T}=\begin{cases}\hat{\mathcal{T}}_{\text{CLS},{m_{1}}}+\hat{\mathcal{T}}_{\text{CLS},{m_{2}}},&m_{1},m_{2}\text{ are available,}\\ \hat{\mathcal{T}}_{\text{CLS},{m_{2}}}+\hat{\mathcal{T}}_{\text{CMPT},{m_{2}}},&m_{1}\text{ is missing,}\\ \hat{\mathcal{T}}_{\text{CLS},{m_{1}}}+\hat{\mathcal{T}}_{\text{CMPT},{m_{1}}},&m_{2}\text{ is missing.}\end{cases}$ (5)
The fused feature token $\mathcal{T}$ is then passed through a linear classifier head to make a prediction $\hat{y}$ as
$\hat{y}=h(\mathcal{T};\Theta_{h}).$ (6)
$\Theta_{h}$ represents the parameters of the classifier head that are learned during training.
Finally, we learn the low-rank factors for each encoder ($(\{\Delta_{f,m}\})$ and classifier head $(\Theta_{h})$ by minimizing the task-specific loss $\mathcal{L}_{\text{task}}$ combined with the alignment loss:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{task}}+\lambda\mathcal{L}_{\text{align}},$ (7)
where $\lambda$ is a hyperparameter that we set to $\lambda=0.20$ in all our experiments.
Our model can handle missing modalities during both training and inference. During training on complete modality data, we apply random modality dropout, where we either use both modalities or randomly drop one in each iteration. This way, the model learns to make correct predictions in missing modality scenarios using the CMPTs from the available modality. When training data contains missing modalities, we learn the CMPTs by optimizing the alignment loss using only the samples with complete modalities, as shown in (4). For missing modalities, we employ placeholders (e.g., an empty string for text and zeros for images, audio, and video) following standard practices [22, 39], and use the CMPT from the available modality as an approximation to make predictions, as described in (5). To enhance training efficiency, we leverage LoRA [15] with a rank of 1 to fine-tune the encoders in a parameter-efficient manner. LoRA layers are inserted after the query, key, value, and output layers in every attention blocks within the encoders. Our experiments demonstrate that LoRA with rank 1 is sufficient to achieve better performance compared to existing state-of-the-art methods, as discussed in Section 4.
4 Experiments and Results
Table 1: Performance comparison when modality gets missing during both training and inference. We train and evaluate our model with same modality ratio following [22, 16, 20] and report the mean and standard deviation (SD) of three independent runs with different seeds.
Datasets % availability ViLT MAP [22] MSP Visual- Ma et al. Kim et al. MuAP [6] CMPT (Ours)
Image Text [21] Input Attn. [16] BERT [23] [29] [20] Head Cross Mean $\pm$ SD
MM-IMDb (F1-Macro) 30% 100% 37.61 46.30 44.74 47.45 38.63 46.63 56.03 47.21 46.73 60.21 $\pm$ 0.40
65% 65% 36.30 42.66 41.56 42.03 37.23 41.28 49.24 42.57 43.92 54.04 $\pm$ 0.23
100% 30% 34.71 39.22 38.16 38.34 36.41 38.65 43.21 41.37 39.88 52.61 $\pm$ 0.12
UPMC Food-101 (Accuracy) 30% 100% 76.93 86.18 86.05 86.34 77.41 86.38 87.11 86.90 86.59 87.48 $\pm$ 0.13
65% 65% 69.03 79.08 78.09 78.89 71.06 78.58 82.67 77.87 78.95 82.83 $\pm$ 0.25
100% 30% 66.29 74.53 72.57 73.77 67.78 73.41 78.81 74.61 74.60 80.37 $\pm$ 0.18
Table 2: Performance comparison when a complete modality gets missing during inference. We report % accuracy for all the datasets. Our method shows overall better performance compared to recent state-of-the-art methods.
Datasets Available Modality Missing Modality FiLM BiGated OGM-GE QMF MLA CMPT
[37] [18] [36] [54] [55] (Ours)
CREMA-D Audio Video 53.89 51.49 53.76 59.41 59.27 67.20
Video Audio 18.67 17.34 28.09 39.11 64.91 76.21
Audio - Video - 60.07 59.21 68.14 63.71 79.70 88.84
KS Audio Video 48.67 49.96 48.87 51.57 54.67 68.27
Video Audio 23.15 23.77 29.73 32.19 51.03 85.77
Audio - Video - 63.33 63.72 65.74 65.78 71.35 91.21
UPMC Food-101 Image Text 4.68 14.20 22.35 45.74 69.60 75.66
Text Image 85.84 85.79 85.17 84.13 86.47 80.66
Image - Text - 87.21 88.87 87.54 92.87 93.33 94.47
4.1 Datasets
We evaluate our approach on five popular multimodal datasets across different tasks. A brief description of each dataset is provided here, with further details in Section S2 in the supplementary.
UPMC Food-101 [46] is a multimodal classification dataset consisting of 101 food classes. It has a total of 90,704 image-text pairs, divided into a training set of 67,988 pairs and a test set of 22,716 pairs.
MM-IMDb dataset [2] contains 25,959 samples with image and text modalities for multi-label movie genre classification. It is split into 15,552 training, 2,608 validation, and 7,799 test samples across 23 classes.
Kinetics-Sound (KS) [1] is a subset of Kinetics400 dataset [17] for human action recognition using audio and video as input modalities. It has 31 classes, with 14,739 samples in the training set and 2,594 samples in the test set.
Audio-Visual Event (AVE) localization dataset [42] is used for multimodal event localization. It includes 4,143 10-second videos divided into train/val/test splits containing 3,339/402/402 videos, respectively. The dataset spans 28 event categories.
CREMA-D dataset [4] includes 7,442 short video clips from 91 actors expressing six emotion categories. It contains 6,698 samples for training and 744 for testing. The dataset is used for multimodal emotion recognition.
4.2 Implementation details
Vision-Language datasets. We use pretrained ViT-B [8] from CLIP model [38] as image encoder and BERT-base-uncased model [27] as text encoder. The max token length is set to 40 for UPMC Food-101 and 256 for MM-IMDb dataset, respectively.
Audio-Video datasets. We use AST model [13] pretrained on AudioSet dataset [12] as audio encoder and ViT-B [8] from CLIP model [38] as video encoder. Raw audio is converted to a 128-bin mel spectrogram at 16 kHz, with a maximum length of 1024 frames. For video, we sample 3 frames randomly while training and uniformly while testing. Other pre-processing steps are similar to [55, 36, 9].
Hyperparameter settings. We use Python 3.8.19 and PyTorch 2.2.2 for training and evaluating our models. All the models are trained using two NVIDIA RTX 2080Ti GPUs. For vision-language datasets, we set the learning rate to $10^{-3}$ and train the models for 10 epochs with a batch size of 8. For audio-video datasets, the learning rate is set to $5\times 10^{-5}$ and models are trained for 100 epochs with a batch size of 4. We use AdamW [26] optimizer with $\epsilon=10^{-8}$ and weight decay = 0.02. Cross entropy loss is applied while training the models. We utilize a polynomial learning rate scheduler with power=0.9 and set the first 5 epoch as warm-up. The learning rate is set to 0.1 times the original learning rate during warm-up epochs. We set LoRA [15] rank = 1 and insert them after query, key, value and output layers of each transformer block. Further details with all the hyperparameters can be found in Section S3 in the supplementary.
For each task and dataset, we compare against state-of-the-art published work. For this reason, the set of baselines varies across experiments based on the availability of published results. In each table, the best and second best results are bold and underlined, respectively.
Table 3: Performance comparison with both modalities available during training and inference. We report % accuracy for UPMC Food-101 and F1-Macro/F1-Micro score for MM-IMDb dataset. ‘-’ indicate results that are not reported in corresponding papers.
Methods Parameters (M) UPMC MM-IMDb
Total Learnable Food-101
ViLT [21] 112.63 112.63 92.00 55.30 / 64.70
MBT [33] 196.00 196.00 93.56 59.60 / 64.81
MMBT [19] 170.00 170.00 94.10 60.80 / 66.10
MMLoRA [9] 196.10 196.10 93.70 61.70 / 67.20
PMF [24] 198.43 2.54 91.51 58.77 / 64.51
PMF-L [24] 643.95 4.44 91.68 61.66 / 66.72
MLA [55] 218.26 218.26 93.33 - / -
CMPT (Ours) 195.51 0.27 94.47 63.58 / 69.23
4.3 Performance when modalities are missing
Existing literature considers missing modalities at both training and inference or only at inference. Our method achieves state-of-the-art performance in both scenarios, as detailed in the next two sections.
4.3.1 Missing modality during training and inference
In the first setup, a modality is missing at a fixed rate during both training and inference, as studied in [22, 16, 20]. We follow the experiment setup in [22, 16, 20] and compare our method against seven state-of-the-art approaches on the MM-IMDb and UPMC Food-101 datasets. As shown in Table 1, our method consistently outperforms all baselines across various missing modality configurations. Specifically, compared to the most recent approach in Kim et al. [20], we achieve a 4.18%–9.40% improvement in F1-macro score on MM-IMDb dataset and a 0.16%–1.56% improvement in accuracy on UPMC Food-101 dataset across different missing modality scenarios. These results also suggest that the CMPTs can provide good proxy for the missing class tokens, which significanctly improves the model robustness. To ensure a reliable evaluation, we run our method with three different random seeds and report the mean and standard deviation (SD) for various missing scenarios. The low SD values indicate the stability and reliability of our method. We also assessed the performance of the models with different distributions of modality ratios at training and testing, for which the results are presented in Section S4 of the supplementary material.
4.3.2 Missing modality during inference only
In the second setup, all modalities are available during training, but one modality is completely missing during inference, as examined in [55, 39]. Following the setup in [55], we evaluate our model when all modalities are available during training, but one modality is completely missing during inference. As shown in Table 2, our method outperforms existing baselines in most cases. For instance, when the audio modality is missing, all methods except MLA [55] experience a significant performance drop on the CREMA-D and Kinetics-Sound (KS) datasets. On the CREMA-D dataset, our method achieves 7.93% and 11.30% improvements over the most recent MLA approach, when video and audio are missing, respectively. Similarly, on the KS dataset, our method outperforms MLA with 13.60% and 34.74% improvements when video and audio are missing, respectively. While MLA performs better when the image modality is missing on the UPMC Food-101 dataset, it suffers a larger performance drop when text is missing. In contrast, our method maintains a balanced performance across different missing modalities. Furthermore, across all datasets, our method achieves the best performance when all modalities are available and demonstrates superior robustness compared to other approaches. These results highlight the effectiveness of our method in handling scenarios where a modality is entirely absent during inference.
Table 4: Performance (accuracy) comparison with both modalities available during training and inference on Audio-Video datasets. denotes results generated using existing code.
Method Parameters (M) KS AVE CREMA-D
Total Learnable
G-Blend [45] 22.36 22.36 62.20 65.50 58.70
FiLM [37] 22.35 22.35 63.33 - 60.07
BiGated [18] 22.35 22.35 63.72 - 59.21
OGM-GE [36] 22.35 22.35 65.74 76.90 68.14
QMF [54] 22.36 22.36 65.78  68.16 63.71
PMR [11] 22.36 22.36 - 66.40 61.80
MLA [55] 22.37 22.37 71.35  64.68 79.70
MBT [19] 172.00 172.00 85.00 77.80 -
MMLoRA [9] 172.19 172.19 91.40 96.20 88.60
CMPT (Ours) 172.16 0.17 91.21 96.77 88.84
(a) Dataset: MM-IMDb
(b) Dataset: UPMC Food-101
Figure 3: Generalization to varying missing rates during inference. Models are trained with 100% image + 100% text and evaluated with 100% image + $x$text following [29, 20]. Our approach demonstrates better generalization, particularly under severe modality loss.
4.4 Performance when all modalities are available
A method designed to handle missing modalities should not compromise performance when all modalities are present. In this subsection, we demonstrate that our approach also achieves comparable or superior performance to state-of-the-art models across five datasets when all modalities are available. These results highlights that incorporating CMPTs does not introduce any drawback when all modalities remain available.
To evaluate this, we train and test our method when all modalities are available during training and inference. Results in Table 3 demonstrate that our proposed method achieves state-of-the-art performance on both the UPMC Food-101 and MM-IMDb datasets. Our method surpasses the most recent MLA approach [55] by 1.14% on UPMC Food-101. Our method achieves 1.88% and 2.03% improvement in F1-macro and F1-micro scores, respectively, over second-best performing MMLoRA [9] on MM-IMDb.
We present the results for audio-video datasets in Table 4, where we compare our method against nine state-of-the-art methods. Our approach consistently achieves comparable or superior performance across all three datasets, while using the fewest number of learnable parameters.
4.5 Generalization to different missing rates
In real-world scenarios, the missing rates for modalities during inference can be unpredictable. Any method designed to handle missing modalities needs to generalize effectively across varying levels of missing modality rates during inference. To evaluate this capability, we trained our models on complete multimodal data and assessed their performance under varying levels of missing modalities during inference. Specifically, we follow the experiment setup similar to [29, 20] where models are trained with 100% image + 100% text data and evaluated with 100% image + $x$text.
As illustrated in Figure 3, our method consistently outperforms all baselines across different missing scenarios, demonstrating superior robustness. We observe a decline in performance of all methods as the amount of available text decreases, but our approach maintains stronger resilience compared to other methods. For instance, on the MM-IMDb dataset, our method achieves a 5.8% higher F1-Macro score than the most recent method by Kim et al. [20] and a 25.6% improvement over ViLT [21] when only 10% of text is available. A similar trend is observed on the UPMC Food-101 dataset, where our method consistently yields superior performance, confirming its robustness to missing modalities at inference.
Note that published work in [29, 20, 21] did not report results for different missing rates of images, which prevents us from reporting a direct comparison in that setting. Nevertheless, our results highlight the effectiveness of CMPTs in handling different missing modality scenarios, making it a more generalizable solution for real-world applications.
(a) Available: Text, Missing: Image.
(b) Available: Image, Missing: Text.
Figure 4: Effectiveness of CMPT in handling missing modalities on MM-IMDb dataset. CMPT significantly improves performance when modalities are severely missing. All models are trained on 100% image + 100% text and tested with varying levels of missing modalities.
(a) Available: Image, Missing: Text.
(b) Available: Text, Missing: Image.
Figure 5: CMPT improves performance across most of the classes when modalities are missing. However, it fails to accurately predict certain classes (e.g., short/documentary) in the absence of text, likely due to the image modality’s insufficient information for those classes. Classes are sorted based on the amount of performance improvement.
4.6 Ablation studies
4.6.1 Effectiveness of CMPT
To evaluate the effectiveness of CMPTs in handling missing modalities, we compare our method against two baselines: (i) a standard model trained without modality dropout or CMPT (baseline) and (ii) a model trained with modality dropout, where modalities are randomly removed during training to improve robustness to missing data.
Figure 4 presents the main results for MM-IMDb dataset, where the baseline model experiences a sharp performance drop as the proportion of missing modalities increases. While modality dropout improves robustness, it fails to fully recover the lost performance. Incorporating CMPT further enhances performance, achieving a 32.06% improvement over the baseline and 11.55% over modality dropout when the image modality is entirely missing as shown in Figure 4(a). A similar trend is observed when text is missing as shown in Figure 4(b). Notably, CMPTs become increasingly effective as the amount of missing data increases, demonstrating its strong ability to mitigate performance degradation in the presence of severe modality loss. Similar trend appears in UPMC Food-101 dataset, which we discuss in Sec. S5 in the supplementary material.
4.6.2 Effects of CMPT on different classes
When analyzing the effectiveness of CMPT, a key question is whether the overall performance gains are driven by just a few classes or if CMPT consistently improves performance across most classes. To address this, we conducted a per-class performance analysis with and without CMPT.
As shown in Figure 5, CMPT improves performance across nearly all classes in both text-missing and image-missing scenarios. Notably, it significantly mitigates performance degradation in modality-sensitive classes. For example, when text is missing, classes like Horror, Animation, and War experience severe drops in performance, but CMPT effectively recovers much of the loss (Figure 5(a)). Similarly, when images are missing, CMPT provides substantial gains in Thriller, Action, and Romance (Figure 5(b)). A slight drop is observed in Short and Documentary classes when text is missing, likely due to insufficient information in the image modality alone for those classes. These results demonstrate that CMPT enhances performance across most classes, leading to a strong overall performance improvement. A similar trend is observed on the UPMC Food-101 dataset (further discussed in Section S6 of the supplementary material).
5 Limitations and Future Directions
Below we list some limitations and potential future directions. First, while our approach shows improved performance with missing modalities over existing methods, missing modalities still cause performance drop. Second, we observe that certain classes are heavily dependent on a specific modality, which leads to significant performance drop in those classes when that specific modality is missing. A natural question arises as how can we design a better training strategy that encourages more uniform reliance on different modalities while maintaining overall performance in both complete and missing modality scenarios? Future work can also explore developing a general framework for explaining modality contribution and expected performance drop under different missing modality scenarios.
6 Conclusion
We presented a simple yet effective framework for utilizing pre-trained unimodal encoders for robust multimodal learning. Our proposed framework enables training of multimodal systems that offer robust performance under different scenarios of missing modalities during training and inference. To compensate for missing modalities, we propose to learn cross-modal proxy tokens (CMPTs) to approximate the missing modality class tokens from the available modality using an alignment loss during training. Instead of training the unimodal encoders from scratch, we only learn rank-1 adapters along with task-specific classifier heads. Our extensive evaluations across diverse tasks and datasets demonstrate the robustness and adaptability of our approach in both complete and missing modality scenarios. Our analysis also reveals that CMPTs can generalize well to different missing scenarios and enhance performance across most of the classes to provide and overall performance boost. Overall, our approach provides a robust and efficient solution for multimodal learning that maintains high performance in different real-world scenarios.
7 Acknowledgment
This work is supported in part by an Amazon Gift award and NSF awards 2046293 and 2406199.
Supplementary Material
S1 Related Work on Multimodal Learning
In this section, we provide a general overview of multimodal learning, which involves integrating information from different input sources to improve performance on downstream tasks. We can broadly categorize multimodal learning approaches into two main categories: designing specialized models and adapting existing models.
Designing specialized models to effectively fuse multiple input sources is a common strategy in multimodal learning. For vision-language tasks, models such as ViLT [21], ViLBERT [27], and Visual-BERT [23] process image patches and text tokens directly within a transformer architecture for tasks like image captioning and visual question answering. MMBT [19] integrates both text and images within a BERT-based framework [7] to model cross-modal dependencies. MBT [33] leverages fusion bottlenecks for efficient modality-specific interaction, particularly in audio-video tasks. Specialized multimodal models have also been proposed for other applications, including action recognition [51, 56], segmentation [53, 49], and sentiment analysis [51, 43]. While these models excel in specific tasks, extending them to new modalities or domains often proves challenging and resource-intensive, as it requires retraining or significant modification of the model architecture.
Adapting existing models offers an alternative approach to multimodal learning, focusing on the use of pre-trained unimodal models and adapting them for multimodal tasks. One such strategy is gradient modulation, where techniques like G-Blend [45] modulate gradients to combine unimodal and multimodal knowledge. OGM-GE [36] improves upon this by emphasizing modality-specific information, enhancing the model’s ability to leverage unique features from each input source. PMR [11] further refines this approach by modulating parameter groups based on the relevance of each modality, while QMF [54] introduces quadratic modulation for adaptive fusion. More recently, MLA [55] proposed alternating adaptation, which toggles between unimodal adaptation and gradient modulation to enable more efficient cross-modal learning. Another notable approach, MMLoRA [9], uses a two-phase strategy: unimodal fine-tuning followed by multimodal adaptation, which helps achieve optimal performance across a variety of tasks. However, while these methods allow for effective cross-modal learning, they often require significant computational resources and learning a large number of parameters to achieve optimal performance on downstream tasks.
Table 5: List of all the hyperparameters used in our experiments.
Hyperparameters Image-Text Audio-Video
Datasets Datasets
Learning Rate $10^{-3}$ $5\times 10^{-5}$
Scheduler Polynomial Polynomial
Power $0.9$ $0.9$
Num Epochs $10$ $100$
Warmup Epochs $5$ $5$
Optimizer AdamW AdamW
$\epsilon$ $10^{-8}$ $10^{-8}$
Weight Decay $0.02$ $0.02$
$\lambda$ $0.20$ $0.20$
LoRA Rank $1$ $1$
LoRA Alpha $1$ $1$
LoRA Dropout $0.1$ $0.1$
Image Encoder ViT-B ViT-B
Text Encoder BERT-Base -
Audio Encoder - AST
Image Resolution $224\times 224$ $224\times 224$
Patch Size $16\times 16$ $16\times 16$
Batch Size 8 4
Audio Sampling Rate - 16 kHz
Frequency Bins - $128$
Maximum Length - 1024 Frames
Frames Per Video - 3
Python Version 3.8.19 3.8.19
PyTorch Version 2.2.2 2.2.2
GPU Model RTX 2080Ti RTX 2080Ti
Num GPUs $2$ $1$
Table 6: Performance comparison when modality gets missing during both training and inference. We train and evaluate our model with different modality ratio following [16, 20] and report the mean and standard deviation (SD) of three independent runs with different seeds.
Datasets Missing Rate ($\eta$) Train Inference ViLT MAP MSP Kim et al. CMPT (Ours)
Image Text Image Text [21] [22] [16] [20] (Mean $\pm$ SD)
MM-IMDb (F1-Macro) 70% 30% 100% 30% 100% 37.63 47.18 47.45 56.03 60.21 $\pm$ 0.40
65% 65% 34.47 37.25 42.17 49.81 50.05 $\pm$ 0.53
100% 30% 30.00 24.31 34.32 43.07 36.10 $\pm$ 0.19
65% 65% 30% 100% 36.48 42.72 44.81 55.26 57.45 $\pm$ 0.21
65% 65% 36.54 40.84 42.03 49.24 54.04 $\pm$ 0.23
100% 30% 33.76 37.80 37.30 42.46 50.05 $\pm$ 0.29
100% 30% 30% 100% 27.25 20.17 34.11 54.67 44.72 $\pm$ 0.46
65% 65% 32.51 29.79 36.97 49.09 49.26 $\pm$ 0.42
100% 30% 34.26 36.89 38.34 43.21 52.61 $\pm$ 0.12
UPMC Food-101 (Accuracy) 70% 30% 100% 30% 100% 76.40 86.34 86.34 87.11 87.48 $\pm$ 0.13
65% 65% 59.03 57.92 71.88 81.22 77.52 $\pm$ 0.29
100% 30% 41.94 29.71 56.87 75.41 67.60 $\pm$ 0.26
65% 65% 30% 100% 73.62 85.55 85.91 87.35 87.36 $\pm$ 0.20
65% 65% 68.60 78.49 78.89 82.67 82.83 $\pm$ 0.25
100% 30% 64.25 71.17 71.58 78.13 78.40 $\pm$ 0.09
100% 30% 30% 100% 43.67 27.46 57.22 86.90 85.34 $\pm$ 0.15
65% 65% 54.75 50.79 65.13 82.35 82.78 $\pm$ 0.22
100% 30% 66.01 73.71 73.77 78.81 80.37 $\pm$ 0.18
S2 Detailed Dataset Description
UPMC Food-101 dataset [46] is a popular and challenging multimodal classification dataset. It has two input modalities: image and text. The dataset has 90,704 samples divided into training and test sets having 67,988 and 22,716 samples, respectively. It contains 101 classes, which are the same as those in the ETHZ Food-101 dataset [3]. The samples are noisy, each category contains around 5% irrelevant images as they were collected in an uncontrolled environment without any human human intervention during the data collection process.
MM-IMDb dataset [2] is a widely used multimodal dataset for multi-label movie genre classification task. It consists of 25,959 samples, each containing both image and text as input modality. The dataset is divided into three subsets: 15,552 samples for training, 2,608 for validation, and 7,799 for testing. It has 23 classes and each sample can have one or more genre. This dataset serves as a standard benchmark for evaluating multimodal models in the multi-label classification task.
Kinetics-Sound (KS) dataset [1] is a subset of the Kinetics-400 dataset [17]. It is used for human action recognition using audio and video as input modalities. This dataset includes 31 different action classes that span a variety of everyday human activities, making it suitable for evaluating multimodal models. The dataset is split into two subsets: a training set containing 14,739 samples and a test set containing 2,594 samples.
Audio-Visual Event (AVE) Localization dataset [42] is a benchmark dataset used for multimodal event localization task. It contains 4,143 10-second videos. The dataset is divided into training, validation, and test sets, with 3,339 videos for training, 402 for validation, and 402 for testing. The dataset encompasses 28 different event categories providing a broad range of scenarios where both audio and visual signals are crucial for accurate event detection. All videos in the AVE dataset are collected from YouTube.
CREMA-D dataset [4] is used for multimodal emotion recognition. It has audio and video as input modalities. The dataset contains 7,442 short video clips performed by 91 actors. The clips cover six emotions: angry, happy, sad, neutral, disgust, and fear. Emotion labels were obtained from 2,443 crowd-sourced raters to ensure diverse evaluations. The dataset is divided into 6,698 samples for training and 744 samples for testing.
(a) Available: Text, Missing: Image.
(b) Available: Image, Missing: Text.
Figure 6: Effectiveness of CMPT in handling missing modalities on UPMC Food-101 dataset. CMPT significantly improves performance when modalities are severely missing. Models are trained on 100% image + 100% text and tested with varying levels of missing modalities.
S3 Implementation Details
In Section 4.2 of the main paper, we briefly discussed the hyperparameter settings used in our experiments. Here, we provide a complete list of all the hyperparameters and their values in Table 5. The “Image-Text Datasets” column includes the hyperparameters for UPMC Food-101 and MM-IMDb datasets, while the “Audio-Video Datasets” column lists the hyperparameters for Kinetics-Sound, AVE, and CREMA-D datasets.
S4 Different Modality Ratios During Training and Inference
A robust multimodal learning method should maintain strong performance even when the missing modality ratio during inference differs from that during training. To evaluate the robustness of our approach, we follow the experimental setup of [16, 20] and train and test our model with varying modality ratios. Let $\eta$ denote the missing rate. We consider three scenarios: missing-text, missing-image, and missing-both. In the first two cases, $\eta\%$ of the samples contain only images or only text, while $(1-\eta)\%$ retain both modalities. For the missing-both case, $\frac{\eta}{2}\%$ of the samples contain only texts, $\frac{\eta}{2}\%$ contain only images, and $(1-\eta)\%$ include both modalities. We set $\eta=70$ throughout our experiments, following [16, 20].
As summarized in Table 6, our method consistently outperforms existing state-of-the-art approaches in most scenarios. While the method proposed by Kim et al. [20] performs well in certain cases, our approach demonstrates superior performance across both the MM-IMDb and UPMC Food-101 datasets. To ensure a rigorous evaluation, we report the mean and standard deviation (SD) over three independent runs with different seeds. The low SD values further indicate the stability and reliability of our method. These results substantiate our claim that CMPTs can effectively estimate missing modality features, enhancing robustness to varying missing modality conditions. Other methods reported in Table 1 do not report results for this specific experimental setup, and thus we are unable to compare with them.
S5 Effectiveness of CMPT
We extend the study from Section 4.6.1 to evaluate the effectiveness of CMPT in handling missing modalities on the UPMC Food-101 dataset. We compare it against two baselines: (i) a standard model trained without modality dropout or CMPT (baseline) and (ii) a model trained with modality dropout, where modalities are randomly removed during training to improve robustness to missing data.
As illustrated in Figure 6, the baseline model suffers a sharp performance decline as the proportion of missing modalities increases. While modality dropout improves robustness, it fails to fully recover the lost performance. Incorporating CMPT further enhances performance, achieving a 5.07% improvement over the baseline and 3.81% over modality dropout when the image modality is entirely missing as shown in Figure 6(a). A similar trend is observed when text is missing as shown in Figure 6(b)), where CMPT achieves 25.59% and 11.96% higher accuracy compared to the baseline and modality dropout, respectively. Consistent with our previous findings, CMPT becomes increasingly effective as the proportion of missing data increases, demonstrating its strong capability to mitigate performance degradation in the presence of severe modality loss.
Figure 7: CMPT improves performance across most of the classes when text modality is missing on UPMC Food-101 dataset. However, it fails to accurately predict certain classes (e.g., bibimbap/sashimi), likely due to the image modality’s insufficient information for those classes. Classes are sorted based on the amount of performance improvement.
Figure 8: CMPT improves performance across most of the classes when image modality is missing on UPMC Food-101 dataset. However, it fails to accurately predict certain classes (e.g., lobster_roll_sandwitch/cheese_plate), likely due to the text modality’s insufficient information for those classes. Classes are sorted based on the amount of performance improvement.
Figure 9: Per class performance analysis on MM-IMDb dataset. Certain classes show strong dependence on a specific modality.
Figure 10: Per class performance analysis on UPMC Food-101 dataset. Certain classes show strong dependence on a specific modality.
S6 Does CMPT Benefit All Classes?
To further evaluate CMPT’s effectiveness, we conduct a per-class performance analysis on the UPMC Food-101 dataset, following the approach from Section 4.6.2.
As shown in Figure 7 and 8, CMPT improves or maintains performance across most of the classes in both text-missing and image-missing scenarios. Notably, it mitigates performance degradation in modality-sensitive classes. For instance, when text is missing, classes such as guacamole, pancakes, and foie_gras experience severe drops in performance, but CMPT effectively recovers much of the loss as shown in Figure 6(b). Similarly, in image-missing scenarios, CMPT provides substantial gains for sushi, cup_cakes, and pizza as shown in Figure 4(a). However, a slight performance drop is observed in certain classes when either text or image is absent, likely due to insufficient information in the available modality alone for those classes. Overall, these results reinforce our previous findings, demonstrating that CMPT enhances performance across most classes, leading to a strong overall performance improvement.
S7 Analysis of Modality Dependence
We observe that the impact of missing modalities varies significantly across classes. Certain classes show strong dependence on a specific modality. In this experiment, we train our model on modality-complete data and evaluate it on both modality-complete and modality-incomplete test samples.
Figure 9 presents per-class performance on the MM-IMDb dataset, revealing heterogeneous effects of missing modalities. For instance, the Drama and Thriller classes maintain consistent performance regardless of missing modalities. In contrast, the Biography and Documentary classes experience significant decline in performance when text is missing, even if images are available. Interestingly, certain classes like Animation show slightly higher performance with single-modality inputs, particularly in image-only scenario. Similar patterns emerge in Sport and Music categories, where text-only samples performs slightly better than complete modality. This discrepancy may stem from the simple sum-fusion strategy we use, which might not effectively model complex modality interactions. These findings indicate that the effect of missing modalities is not uniform across classes, with some being more modality-dependent than others.
Extending this analysis to UPMC Food-101, we observe that multimodal (image-text) data consistently outperforms unimodal inputs across all classes as shown in Figure 10. However, the reliance on specific modalities varies widely among food categories.
Classes such as dumplings, Frozen_yogurt, and Ice cream depend heavily on text, showing strong performance with text-only inputs but experiencing significant drops when text is missing. Conversely, categories like tuna_tartare, beef_carpaccio, and beet_salad rely more on image information, with performance deteriorating notably when images are absent. This observation suggests that some classes rely more on texts, while others depend heavily on images for accurate classification.
Interestingly, some classes such as beef_tartare, chicken_quesadilla, and croque_madame exhibit poor performance with unimodal inputs but improve significantly with both image and text, reinforcing the importance of multimodal data. Performance on these classes improve significantly when we utilize both image and text. Overall, these findings confirm that while multimodal information universally enhances performance, the degree of modality dependence remains heterogeneous across classes.